# Sharper Guarantees for Learning Neural Network Classifiers with Gradient Methods

### v1

@@ E.6 Term I:
We first consider $k\neq 1,2$. By applying Hoeffding’s concentration inequality we obtain $\sum_{j=1}^{n}x_{j}(\ell)x_{j}(k)x_{j}(1)x_{j}(2)<\sqrt{n}\log(dT)$ uniformly for every $\ell\in[d],t\leq T$ w.p. $1-\exp(\log(dT)-2\log^{2}(dT))=:1-p_{x}.$ Applying Hoeffding again yields,
$\nabla\hat{F}_{i}^{k}(w^{t})-\nabla\hat{F}_{i}^{k}(\mathbb{E}_{x}[w^{t}])$ $=\frac{a_{i}}{nm}\sum_{j=1}^{n}\langle x_{j},\mathbb{E}_{x}[w_{i}^{t}]-w_{i}^{t}\rangle x_{j}(k)x_{j}(1)x_{j}(2)$
@@
$\Pr_{x^{t}}\left(\forall i,k,t:\nabla\hat{F}_{i}^{k}(w^{t})-\nabla\hat{F}_{i}^{k}(\mathbb{E}_{x}[w^{t}])\leq\frac{\|v^{t}\|}{\sqrt{n}m}\log(dT)\log(dmT)\Big{|}v^{t}\right)\geq 1-p_{2}-p_{x}.$
@@
$\Pr_{x^{t},v^{t}}\left(\forall i,k,t:\nabla\hat{F}_{i}^{k}(w^{t})-\nabla\hat{F}_{i}^{k}(\mathbb{E}_{x}[w^{t}])\leq\frac{\|v^{t}\|}{\sqrt{n}m}\log(dT)\log(dmT)\right)\geq 1-p_{2}-p_{x}.$
@@
$\nabla\hat{F}_{i}^{k}(w^{t})-\nabla\hat{F}_{i}^{k}(\mathbb{E}_{x}[w^{t}])$ $=\frac{a_{i}}{nm}\sum_{j=1}^{n}\langle x_{j},\mathbb{E}_{x}[w_{i}^{t}]-w_{i}^{t}\rangle x_{j}(2)$
@@
$\nabla\hat{F}_{i}^{k}(w^{t})-\nabla\hat{F}_{i}^{k}(\mathbb{E}_{x}[w^{t}])$ $=\frac{a_{i}}{nm}\sum_{j=1}^{n}\langle x_{j},\mathbb{E}_{x}[w_{i}^{t}]-w_{i}^{t}\rangle x_{j}(1)$
@@ E.7 Combining two terms
We next show that $\|v^{t}(3:d)\|$ is much smaller. To see this, we replace our derived bound in the third equation above. define $\beta^{t}:=\|v^{t}(3:d)\|$ and note that $\beta^{t}\leq\sqrt{d}|v^{t}(k)|$ for every $k\in[3,d]$ and thus:
@@
and by noting that $|P|=|N|=m/2$ for some $m\geq 2$ and selecting $u=m/16$ in the second inequality we have with probability at least $1-e^{-\frac{m}{16}}$,
@@
For the theorem’s choice of $T=\log(d),m=O(\mathrm{poly}(d))$ it is deduced that $e^{\log(m)-\log^{2}(d)}+C(e^{-\log^{2}(d)}+e^{\log(dmT)-2\log^{2}(dmT)}+e^{\log(dT)-2\log^{2}(dT)}+e^{-\frac{1}{2}\log^{2}(dm)+\log(dm)})=o_{d}(1)$. Moreover, for the theorem’s choice $c=7$ and large enough $d$ it holds that $d-4Cd/\log^{\frac{c-6}{2}(d)}\geq 0.$ This leads to the following lower bound on test accuracy which holds with probability at least $1-2e^{-\frac{m}{16}}-e^{\log(m)-\log^{2}(d)}-o_{d}(1)$ over the randomness in data sampling and initialization,
$\Pr_{x,y}\left(y\Phi(w^{t},x)\geq 0\right)\geq 1-\frac{1}{d},$
This completes the proof of Theorem 2.4.

### v2

@@ E.6 Term I:
We first consider $k\neq 1,2$. By applying Hoeffding’s concentration inequality we obtain
$\sum_{j=1}^{n}x_{j}(\ell)x_{j}(k)x_{j}(1)x_{j}(2)<\sqrt{n}\log(dT)$ uniformly for every $\ell\in[d],t\leq T$ w.p. $1-\exp(\log(dT)-2\log^{2}(dT))=:1-p_{x}.$ Applying Hoeffding again yields,
$\nabla F_{i}^{k}(w^{t})-\nabla F_{i}^{k}(\mathbb{E}_{x}[w^{t}])$ $=\frac{a_{i}}{nm}\sum_{j=1}^{n}\langle x_{j},\mathbb{E}_{x}[w_{i}^{t}]-w_{i}^{t}\rangle x_{j}(k)x_{j}(1)x_{j}(2)$
@@
$\Pr_{x^{t}}\left(\forall i,k,t:\nabla F_{i}^{k}(w^{t})-\nabla F_{i}^{k}(\mathbb{E}_{x}[w^{t}])\leq\frac{\|v^{t}\|}{\sqrt{n}m}\log(dT)\log(dmT)\Big{|}v^{t}\right)\geq 1-p_{2}-p_{x}.$
@@
$\Pr_{x^{t},v^{t}}\left(\forall i,k,t:\nabla F_{i}^{k}(w^{t})-\nabla F_{i}^{k}(\mathbb{E}_{x}[w^{t}])\leq\frac{\|v^{t}\|}{\sqrt{n}m}\log(dT)\log(dmT)\right)\geq 1-p_{2}-p_{x}.$
@@
$\nabla F_{i}^{k}(w^{t})-\nabla F_{i}^{k}(\mathbb{E}_{x}[w^{t}])$ $=\frac{a_{i}}{nm}\sum_{j=1}^{n}\langle x_{j},\mathbb{E}_{x}[w_{i}^{t}]-w_{i}^{t}\rangle x_{j}(2)$
@@
$\nabla F_{i}^{k}(w^{t})-\nabla F_{i}^{k}(\mathbb{E}_{x}[w^{t}])$ $=\frac{a_{i}}{nm}\sum_{j=1}^{n}\langle x_{j},\mathbb{E}_{x}[w_{i}^{t}]-w_{i}^{t}\rangle x_{j}(1)$
@@ E.7 Combining two terms
We next show that $\|v^{t}(3:d)\|$ is much smaller. To see this, we replace our derived bound in the third equation above. define $\beta^{t}:=\|v^{t}(3:d)\|$ and note that $\beta^{t}\leq\sqrt{d}|v^{t}(k)|$ for some $k\in[3,d]$. Thus:
@@
and by noting that $|P|=|N|=m/2$ and selecting $u=m/16$ in the second inequality we have w.p. at least $1-e^{-\frac{m}{16}}$,
@@
For the theorem’s choice of $T=\log(d),m=O(\mathrm{poly}(d))$ it is deduced that $e^{\log(m)-\log^{2}(d)}+C(e^{-\log^{2}(d)}+e^{\log(dmT)-2\log^{2}(dmT)}+e^{\log(dT)-2\log^{2}(dT)}+e^{-\frac{1}{2}\log^{2}(dm)+\log(dm)})=o_{d}(1)$. Moreover, for the theorem’s choice $c=7$ and large enough $d$ it holds that $d-4Cd/\log^{\frac{c-6}{2}(d)}\geq 0.$ This leads to the following lower bound on test accuracy,
$\Pr_{x,y}\left(y\Phi(w^{t},x)\geq 0\right)\geq 1-\frac{1}{d},$ (57)
with probability at least $1-e^{-\frac{m}{16}}-e^{\log(m)-\log^{2}(d)}-o_{d}(1)$ over data sampling and initialization. This completes the proof of Theorem 2.4.